Mary Jo wants to buy a boat that is available at two dealerships. The price of the boat is the same at both dealerships. Middlefield Motors would let her make quarterly payments of 3,657.97 dollars for 12 years at a quarterly interest rate of 3.18 percent. Her first payment to Middlefield Motors would be due immediately. If Fairfax Boats would let her make equal monthly payments of 3,027.97 for 3 years and if her first payment to Fairfax Boats would be in 1 month, then what is the monthly interest rate that Mary Jo would be charged by Fairfax Boats? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1
Middlefield Motors
Rate Quarterly interest rate 3.18%
Nper Number of quarterly payment 48 (12*4)
Pmt Quarterly Payment $3,657.97
Type Payment at beginning of quarters 1
PV Present Value of Payments $92,275.13 (Using PV function of excel with Rate=3.18%,Nper=48,Pmt=-3657.97,Type=1)
Price of boat $92,275.13
Fairfax Boats
Pmt Monthly Payment $3,027.97
Nper Number of payments=3*12 36
Pv Price of boat $92,275.13
Type Payment at end of months 0
RATE Monthly Interest Rate 0.93% (Using RATE function of excel with Nper=36,Pmt=3027.97,Pv=-92275.13Type=0)
Monthly Interest Rate          0.0093

Lipuvaru வாயாடாட D7 X fx =PV(D3,D4,-D5,,1) DE E В F G H I J K L 2 3 Rate 4 Nper 5 Pmt 6 Type 7 PV Middlefield Motors QuarterlD15 x fc =RATE(D12,011,-D13,,0) D E F G H к 2 3 Rate 4 Nper 5 Pmt 6 Type 7 PV Middlefield Motors Quarterly interest rate Numb